OpenStudy (anonymous):

Solve the equation x^2 + 5x = 0. A) x = 0, –5 B) x = 0, 5 C) x = 1, –5 D) x = 1, 5

OpenStudy (luigi0210):

Try factoring out an x, what are you left with?

OpenStudy (anonymous):

x^2 + 5x = 0. taking common x(x+5)=0 x= 0 x= -5

OpenStudy (anonymous):

Cookies *~*

OpenStudy (anonymous):

i think its 1.5

OpenStudy (luigi0210):

Annas you noob.

OpenStudy (anonymous):

quadratics are of the form ax^2 + bx + c and can be factored into two binomials (x + f)(x + g). in this equation c = 0 so that means f or g must be 0; e.g., one solution is 0. x(x + g) = ax^2 + bx + c x(x + g) = x^2 + 5x x(x + 5) = x^2 + 5x x(x + 5) = 0 by the zero product property x + 5 has to be 0 or x = 0, so solve x + 5 = 0 to get your answer

OpenStudy (anonymous):

B) x = 0, 5

OpenStudy (anonymous):

ok thank you @Luigi0210 @annas @lilyryan

OpenStudy (anonymous):

Luigi you monkey *~* welcome Tina :3